# SoTA with Less: MCTS-Guided Sample Selection for Data-Efficient Visual Reasoning Self-Improvement

### v1

@@ Abstract
In this paper, we present an effective method to enhance visual reasoning with significantly fewer training samples, relying purely on self-improvement with no knowledge distillation. Our key insight is that the difficulty of training data during reinforcement fine-tuning (RFT) is critical. Appropriately challenging samples can substantially boost reasoning capabilities even when the dataset is small. Despite being intuitive, the main challenge remains in accurately quantifying sample difficulty to enable effective data filtering. To this end, we propose a novel way of repurposing Monte Carlo Tree Search (MCTS) to achieve that. Starting from our curated 70k open-source training samples, we introduce an MCTS-based selection method that quantifies sample difficulty based on the number of iterations required by the VLMs to solve each problem. This explicit step-by-step reasoning in MCTS enforces the model to think longer and better identifies samples that are genuinely challenging. We filter and retain 11k samples to perform RFT on Qwen2.5-VL-7B-Instruct, resulting in our final model, ThinkLite-VL. Evaluation results on eight benchmarks show that ThinkLite-VL improves the average performance of Qwen2.5-VL-7B-Instruct by 7%, using only 11k training samples with no knowledge distillation. This significantly outperforms all existing 7B-level reasoning VLMs, and our fairly comparable baselines that use classic selection methods such as accuracy-based filtering. Notably, on MathVista, ThinkLite-VL-7B achieves the SoTA accuracy of 75.1, surpassing Qwen2.5-VL-72B, GPT-4o, and O1. Our code, data, and model are available at https://github.com/si0wang/ThinkLite-VL.
Figure 1: Recent “Reasoning VLMs” studies finetune “Base VLMs” with extra reasoning training data to improve visual reasoning. This paper presents a data-efficient self-improving method for better training reasoning VLMs. (Left) Comparison of VLMs with different parameter sizes on MathVista. Our model ThinkLite-VL-7B achieves the state-of-the-art (SoTA) accuracy of 75.1, surpassing Qwen2.5-VL-72B-Instruct, GPT-4o, O1, and other 7B-level reasoning VLMs. (Right) Comparison of the reasoning training data size used by 7B-level reasoning models. Our model achieves SoTA performance using only 11k data, and without any additional knowledge distillation.
@@ 1 Introduction
Leveraging long chain-of-thought reasoning with effective reflection during inference, large language models (LLMs) [24, 34] are capable of solving complex reasoning tasks such as math and coding. Recent studies [16] show that large-scale reinforcement fine-tuning (RFT) is a critical factor in enhancing model’s reasoning performance. Notably, substantial reasoning performance improvements can be achieved solely through reinforcement fine-tuning in the post-training stage, even without the standard supervised fine-tuning (SFT) in post-training.
@@
Figure 2: Performance comparison on 8 visual benchmarks. Our model significantly outperforms Qwen2.5-VL-7b-Instruct and other 7b-level reasoning models.
Despite the notable successes in enhancing LLM reasoning with large-scale RFT, similar progress in vision-language models (VLMs) remains limited, likely due to the mismatch between the text-focused pre-training and the multimodal nature of VLM post-training tasks. Recent attempts [22, 12, 53, 81] have employed knowledge-distillation via supervised fine-tuning before the RFT stage, to encourage more visual reasoning related responses being generated. Despite the performance improvement, the knowledge distillation stage is cumbersome, and inherently prevents base VLMs from self-improving themselves in achieving stronger intelligence.
In this paper, we demonstrate that high-quality, appropriately challenging training data is key factor to enable and self-improve visual reasoning ability. When visual reasoning training data aligns properly with the base VLM’s skill level, large-scale RFT alone can significantly enhance VLM’s reasoning ability without relying on knowledge distillation for format fine-tuning or base capability enhancement. Based on this insight, We introduce a data-efficient training pipeline that results in ThinkLite-VL, a reasoning VLM that achieves SoTA visual reasoning performance with less training samples.
The critical factor to ThinkLite-VL’s success is the strategic selection of training samples with suitable difficulty. To achieve this, we repurpose Monte Carlo tree search (MCTS), a classic inference-time search algorithm, to accurately quantify the sample difficulty. Specifically, MCTS’s explicit tree search enforces sufficient thinking compute in deciding the question difficulty, and provide a tight correlation between the question difficulty and the number of MCTS iterations needed to solve it. Our training pipeline begins with collecting 70k open-source samples from three key domains: mathematical reasoning, natural image understanding, and chart comprehension. We then implement MCTS-guided sample selection by applying the VLM itself to perform iterative reasoning on each of the 70k samples, using the number of iterations required to reach the correct solution as a difficulty measure. This rigorous filtering process results in a set of 11k challenging and high-quality samples tailored specifically for our base model. We then directly perform RFT with these selected samples, avoiding any additional supervised fine-tuning steps.
Using the Qwen2.5-VL-7B-Instruct model as our base, we develop our final model, ThinkLite-VL-7B. We evaluate ThinkLite-VL-7B on eight widely used VLM benchmarks. As shown in Figure 2, after RFT with the filtered 11k high-quality data, ThinkLite-VL-7B significantly improves the average performance of Qwen2.5-VL-7B-Instruct from 59.69 to 63.89. It also surpasses the fairly comparable baseline that RFT with the same amount of unfiltered data, from 60.89 to 63.89. Furthermore, compared with the most recent 7B-level reasoning VLMs, ThinkLite-VL-7B consistently demonstrates substantial performance advantages. Notably, on the MathVista benchmark, ThinkLite-VL-7B achieves a state-of-the-art (SoTA) accuracy of 75.1 as shown in Figure 1, significantly surpassing other 7B-level models, open-sourced larger models, GPT-4o, and O1.
@@ 2 Related work
Simulating human-like thinking processes through intermediate reasoning steps has significantly improved the performance of large language models (LLMs) on tasks that require reasoning [24]. One family of methods focuses on explicitly controlling the structure or format of the model’s outputs, such as by applying Chain-of-Thought (CoT) prompting [75] and Self-Consistency [74]. Related lines of work include more elaborate reasoning strategies like Tree of Thoughts [84] or Graph of Thoughts [4]. Additionally, some approaches involve supervised fine-tuning (SFT) on curated datasets with reasoning annotations [50, 85]. Researchers have also explored process reward models (PRMs) that encourage systematic thought processes [33, 64, 68, 27, 94, 45]. Others incorporate search techniques, including Monte Carlo Tree Search (MCTS) or beam search, to refine or verify reasoning paths [77, 78, 5, 15, 19, 70]. Recently, large-scale RL with outcome-based reward functions has been leveraged [16] to elicit powerful reasoning capabilities in LLMs. In this paper, we focus on how to use large-scale RL to enhance the reasoning ability of VLMs.
@@
Vision language models [1, 66, 36, 23, 35, 3, 10, 62, 29, 82] can perform vision tasks using language given visual input through vision encoders like [55, 90, 63]. These models demonstrate comprehensive multimodal capabilities across various scenarios [89, 38, 87, 47, 18, 88, 20, 30] and exhibit reasoning capabilities to some extent [41, 73, 39, 92, 67]. Inspired by the success of reasoning in LLMs, researchers have sought to improve the reasoning capabilities of VLMs. For instance, CoT prompting is applied to VLMs [93, 49, 44, 11, 96, 21] and some papers create multimodal datasets [83, 80, 59, 95, 12, 22, 17, 61], using SFT for knowledge distillation to improve reasoning abilities. Some prior works have also explored improving VLM performance through self-improvement strategies [98, 69, 72, 13]. More recently, RL training has emerged as a promising approach to further strengthen the reasoning capabilities of VLMs [12, 22, 48, 79]. While recent works explore SFT and RL  [12, 22] for VLM reasoning, efficiently utilizing training data and avoiding costly knowledge distillation remains a challenge. In this paper, we propose a novel approach using MCTS to filter for high-quality training instances based on the difficulty level. We then directly apply RL training to enhance reasoning on this curated data, demonstrating strong performance without requiring any SFT stage.
@@
Data filtration aims to identify and retain high-quality, diverse, and task-relevant data while discarding noisy or redundant information to optimize training efficiency and generalization performance. It is important for the pretraining phase [14, 28, 76, 56, 52, 2, 91, 65, 54] and instruction tuning phase [32, 31, 7, 9, 36, 99, 86] of both LLMs and VLMs. In this paper, we specifically focus on filtering training instances to curate data optimally for efficient downstream RL training to improve the reasoning capabilities of VLMs. A concurrent work, MM-Eureka [48], also investigates the impact of data filtration on RFT. However, their approach is limited to a relatively simple self-consistency-based difficulty filtering strategy, where all samples with zero accuracy are discarded. In contrast, we propose a more principled method—MCTS-based sample selection—which enables the identification of truly challenging examples for the VLM. Importantly, our findings reveal that the unsolved samples, which VLMs fail to solve during MCTS, play a critical role in enhancing reasoning performance during RFT, rather than being excluded from the training process.
@@ 3.1 Data Collection
Figure 3: Data statistic of ThinkLite-VL-70k training dataset. We find that converting all answers to open-ended format is critical in reliably assessing question difficulty and effective model training.
@@ 3.2 MCTS-based Sample Selection
Figure 4: Data difficulty distribution of our 11k training set after MCTS-based data filtration. Unsolved refers to data that VLM cannot solve after 50 MCTS iterations.
@@
Specifically, we define the state at step $t$, denoted as $s_{t}$, to represent the prefix of the reasoning chain. The introduction of a new reasoning step, $a$, transitions the state to $s_{t+1}$, which is formed by concatenating $s_{t}$ with $a$. By leveraging VLM itself as policy model, $\pi_{\theta}$, we sample candidate steps from the probability distribution $\pi_{\theta}(a|x,I,s_{t})$, where $x$ denotes the task’s input prompt and $I$ represents the input image. The MCTS process starts from the root node, $s_{0}$, representing the beginning of a sentence. It then iteratively proceeds through three key phases—selection, expansion and simulation—which are described in detail in the subsequent paragraphs. In contrast to previous studies, during the data filtering stage with MCTS, we prioritize computational efficiency and comprehensive exploration of the solution space, with our focus centered on self-rewarding setting. Consequently, throughout the MCTS process, we do not employ any pretrained or separately trained process reward models, thereby simplifying and accelerating the procedure. The prompt used for MCTS is shown in Table 1.
[…]
@@
After selecting a node , we directly utilize the policy $\pi_{\theta}$ to generate several reasoning steps until a final answer is produced or a preset reasoning step limit is reached. Subsequently, we employ the corresponding LLM (in our experiments, the Qwen2.5-VL-7B-Instruct is used, with Qwen2.5-7B-Instruct serving as the critic model) to compare the generated final answer with the ground truth answer, thereby determining the correctness of the response. If the answer is correct, the MCTS process is terminated and the current iteration number $K$ is recorded; if the answer is incorrect, the visit count $N$ of the selected node is updated and the next iteration commences. Table 2 illustrates the prompt employed for the critic model.
@@
We apply this MCTS procedure to the entire collection of 70k data samples and record the iteration number $K$ required to solve each problem, using Qwen2.5-VL-7B-Instruct as the policy model. In this process, $K$ served as a metric for assessing the difficulty of each sample: a higher $K$ indicates that the VLM requires more extensive exploration to arrive at the correct answer, thereby reflecting a greater level of challenge. Ultimately, we select all samples with $K$ greater than 5, as well as those that remained unsolved after 50 iterations, resulting in a final training set of 11k samples. The data difficulty distribution of this final training set is shown in Figure 4.
Table 2: Critic prompt for MCTS simulation results evaluation.

### v2

@@ Abstract
We introduce ThinkLite-VL, a family of visual reasoning models that achieve state-of-the-art (SoTA) performance using an order of magnitude fewer training samples, relying purely on reinforcement fine-tuning (RFT) self-improvement without any knowledge distillation. Our central insight is that sample difficulty critically influences RFT effectiveness: appropriately challenging examples can drive substantial reasoning improvements, even in low-data regimes. However, quantifying sample difficulty in a reliable and scalable manner remains non-trivial. To address this, we repurpose Monte Carlo Tree Search (MCTS) to measure sample difficulty via the number of reasoning iterations a vision-language model (VLM) requires to solve each instance. This MCTS-based selection procedure identifies samples that induce deeper reasoning while remaining solvable, allowing us to filter a high-quality subset from 70k open-source examples spanning math, natural image understanding, and chart comprehension. Using this approach, we select just 11k challenging samples for RFT on Qwen2.5-VL-7B-Instruct and 7.5k samples for Qwen2.5-VL-72B-Instruct. The resulting models, ThinkLite-VL-7B and ThinkLite-VL-72B, significantly outperform their respective base models across eight visual reasoning benchmarks. In particular, ThinkLite-VL-7B improves the average performance of Qwen2.5-VL-7B-Instruct by 7% and surpasses all existing 7B-level models, as well as much larger models such as GPT-4o, O1 and Qwen2.5-VL-72B, achieving a new SoTA score of 75.1 on MathVista. ThinkLite-VL-72B further advances the SoTA frontier, achieving an accuracy of 79.7 on MathVista and an average benchmark improvement of 4.42 over the open-source SOTA. These results demonstrate that MCTS-guided difficulty filtering provides a scalable and effective path toward data-efficient self-improvement in multimodal reasoning. Our code, data, and model are available at https://github.com/si0wang/ThinkLite-VL.
Figure 1: Recent “Reasoning VLMs” studies finetune “Base VLMs” with extra reasoning training data to improve visual reasoning. This paper presents a data-efficient self-improving method for better training reasoning VLMs. (Left) Comparison of VLMs with different parameter sizes on MathVista. Our model ThinkLite-VL-7B achieves the state-of-the-art (SoTA) accuracy of 75.1, surpassing Qwen2.5-VL-72B-Instruct, GPT-4o, O1, and other 7B-level reasoning VLMs. ThinkLite-VL-72B further pushes this boundary to 79.7. (Right) Comparison of the reasoning training data size used by 7B-level and 72B-level reasoning models. Our model achieves SoTA performance using only 11k data (7B) and 7.5k data (72B), and without any additional knowledge distillation.
@@ 1 Introduction
Large language models (LLMs) have demonstrated strong capabilities in solving complex reasoning tasks—such as mathematics and coding—by leveraging chain-of-thought prompting and reflection mechanisms [24, 36]. Recent work [16] highlights the critical role of reinforcement fine-tuning (RFT) in further enhancing reasoning performance. Remarkably, these improvements can be achieved purely via RFT, even without post-training supervised fine-tuning (SFT).
@@
Figure 2: Performance comparison on 8 visual benchmarks. Our model significantly outperforms Qwen2.5-VL-7B and other reasoning models.
However, despite the success of RFT in LLMs, its impact on vision-language models (VLMs) has been less pronounced. A likely cause is the inherent modality gap: VLMs are pretrained on text-heavy objectives, while post-training tasks demand multimodal reasoning. Recent efforts [22, 12, 55, 84] have addressed this by incorporating knowledge distillation and supervised format alignment before RFT. While effective, these pipelines are cumbersome, and fundamentally limit the capacity for models to improve via self-training alone.
In this work, we demonstrate that high-quality and appropriately challenging training samples alone are sufficient to enable self-improvement in VLMs via RFT—without any knowledge distillation. When the training data matches the base model’s capability level, RFT can explore informative rollouts by itself and substantially elevate multimodal reasoning ability. Based on this insight, we introduce ThinkLite-VL, a family of data-efficient reasoning VLMs trained via RFT on a small subset of difficulty-curated examples.
The key to ThinkLite-VL’s performance lies in effective sample selection. We propose to repurpose Monte Carlo Tree Search (MCTS)—a classic inference-time search algorithm—to estimate the difficulty of each training instance. Specifically, we define difficulty as the number of MCTS reasoning iterations a VLM requires to solve a task. This search-based signal tightly correlates with sample difficulty and naturally identifies examples that promote deeper reasoning during training.
Our pipeline begins with 70k open-source samples spanning three core domains: mathematical reasoning, natural image understanding, and chart interpretation. For each example, we simulate an MCTS-based inference trace using the base VLM, and rank samples by the number of reasoning steps required to reach a correct solution. From this pool, we extract two difficulty-filtered subsets: 11k samples for Qwen2.5-VL-7B-Instruct and 7.5k samples for Qwen2.5-VL-72B-Instruct. We then apply RFT directly on these subsets—no supervised fine-tuning or distillation required.
We evaluate our resulting models, ThinkLite-VL-7B and ThinkLite-VL-72B, on eight established VLM benchmarks. After RFT, ThinkLite-VL-7B improves the average performance of Qwen2.5-VL-7B-Instruct from 59.69% to 64.18%, and outperforms a comparable baseline trained on randomly selected 11k samples (60.89%). Similarly, ThinkLite-VL-72B raises the average accuracy of Qwen2.5-VL-72B-Instruct from 68.25% to 72.67%, exceeding the baseline trained on randomly selected 7.5k samples 69.91%.
Furthermore, compared with the most recent 7B-level reasoning VLMs, ThinkLite-VL-7B consistently demonstrates substantial performance advantages as shown in Figure 2. ThinkLite-VL-7B also outperforms much larger models—including GPT-4o, Qwen2.5-VL-72B, and o1—on the MathVista benchmark, achieving a new SoTA score of 75.1% (Figure 1). ThinkLite-VL-72B further advances the frontier, attaining a SoTA accuracy of 79.7% on MathVista.
Our key contributions are:
  1. (1)
    Difficulty as a learning signal. We identify sample difficulty as a critical yet underutilized signal for enabling effective self-improvement in VLMs via RFT, and show the importance of scaling compute for identifying the appropriately challenging training sample.
  2. (2)
    MCTS-guided filtering. We propose a novel use of Monte Carlo Tree Search to estimate sample difficulty by measuring model reasoning iteration count. Across diverse online and offline baselines, MCTS-guided filtering delivers superior performance, benefiting from the explicit tree search.
  3. (3)
    Data-efficient RFT pipeline. We introduce ThinkLite-VL, a data-efficient visual reasoning framework that achieves SoTA performance using only 11k (7B) and 7.5k (72B) training samples, without any knowledge distillation.
  4. (4)
    Strong empirical gains. We demonstrate that ThinkLite-VL-7B and ThinkLite-VL-72B outperform strong baselines and existing SoTA models across eight VLM benchmarks. Notably, ThinkLite-VL-7B improves the average performance of its base model by 7%, and achieves a new SoTA score of 75.1% on MathVista—surpassing larger models such as GPT-4o, O1 and Qwen2.5-VL-72B. ThinkLite-VL-72B further advances this with a MathVista score of 79.7%.
  5. (5)
    Open-source release. We will release the full ThinkLite-VL model family, including both ThinkLite-VL-7B and ThinkLite-VL-72B, and MCTS-filtered training sets for both Qwen2.5-VL-7B and Qwen2.5-VL-72B to support future research in multimodal reasoning.
@@ 2 Related work
Simulating human-like thinking processes through intermediate reasoning steps has significantly improved the performance of large language models (LLMs) on tasks that require reasoning [24]. One family of methods focuses on explicitly controlling the structure or format of the model’s outputs, such as by applying Chain-of-Thought (CoT) prompting [78] and Self-Consistency [77]. Related lines of work include more elaborate reasoning strategies like Tree of Thoughts [87] or Graph of Thoughts [4]. Additionally, some approaches involve supervised fine-tuning (SFT) on curated datasets with reasoning annotations [52, 88]. Researchers have also explored process reward models (PRMs) that encourage systematic thought processes [35, 66, 70, 28, 98, 47]. Others incorporate search techniques, including Monte Carlo Tree Search (MCTS) or beam search, to refine or verify reasoning paths [80, 81, 5, 15, 19, 73]. Recently, large-scale RL with outcome-based reward functions has been leveraged [16] to elicit powerful reasoning capabilities in LLMs. Unlike prior uses of MCTS at inference time [80, 81, 15], we employ MCTS during training to assess sample difficulty and curate a high-impact training subset for RFT. We focus on how to use large-scale RL to enhance the reasoning ability of VLMs.
@@
Vision language models [1, 68, 38, 23, 37, 3, 10, 64, 31, 85] can perform vision tasks using language given visual input through vision encoders like [57, 94, 65]. These models demonstrate comprehensive multimodal capabilities across various scenarios [93, 39, 91, 49, 18, 92, 20, 32] and exhibit reasoning capabilities to some extent [43, 76, 41, 96, 69]. Inspired by the success of reasoning in LLMs, researchers have sought to improve the reasoning capabilities of VLMs. For instance, CoT prompting is applied to VLMs [97, 51, 46, 11, 100, 21] and some papers create multimodal datasets [86, 83, 61, 99, 12, 22, 17, 63], using SFT for knowledge distillation to improve reasoning abilities. Some prior works have also explored improving VLM performance through self-improvement strategies [102, 72, 75, 13]. More recently, RL training has emerged as a promising approach to further strengthen the reasoning capabilities of VLMs [12, 22, 50, 82]. While recent works explore SFT and RL  [12, 22] for VLM reasoning, efficiently utilizing training data and avoiding costly knowledge distillation remains a challenge. In contrast, ThinkLite-VL eliminates the need for SFT or distillation entirely and achieves SoTA performance using just 11k (7B) and 7.5k (72B) samples—an order of magnitude less than prior work. Specifically, we propose a novel approach using MCTS to filter for high-quality training instances based on the difficulty level. We then directly apply RL training to enhance reasoning on this curated data, demonstrating strong performance without requiring any SFT stage.
@@
Data filtration aims to identify and retain high-quality, diverse, and task-relevant data while discarding noisy or redundant information to optimize training efficiency and generalization performance. It is important for the pretraining phase [14, 29, 79, 58, 54, 2, 95, 67, 56] and instruction tuning phase [34, 33, 7, 9, 38, 103, 89] of both LLMs and VLMs. In this paper, we specifically focus on filtering training instances to curate data optimally for efficient downstream RL training to improve the reasoning capabilities of VLMs. A concurrent work, MM-Eureka [50], also investigates the impact of data filtration on RFT. While MM-Eureka [50] filters samples based on zero-shot accuracy, our MCTS-based method provides a more expressive and fine-grained estimate of sample difficulty, capturing both solved and unsolved-but-informative cases. Importantly, our findings reveal that samples requiring extended reasoning—even when not solved by the model—can be highly beneficial during RFT.
To our knowledge, ThinkLite-VL is the first framework to combine search-based sample difficulty estimation with reinforcement fine-tuning—achieving data-efficient self-improvement for visual reasoning at both 7B and 72B scale, without any SFT or distillation.
@@ 3.1 Data Collection
Figure 3: Data statistic of ThinkLite-VL-70k training dataset. We find that converting answers to open-ended format is critical in reliably assessing question difficulty and effective model training.
@@ 3.2 MCTS-based Sample Selection
Figure 4: Data difficulty distribution of our 11k training set after 7B MCTS-based data filtration. Unsolved refers to data that VLM cannot solve after 50 MCTS iterations.
@@
Specifically, we define the state at step $t$, denoted as $s_{t}$, to represent the prefix of the reasoning chain. The introduction of a new reasoning step, $a$, transitions the state to $s_{t+1}$, which is formed by concatenating $s_{t}$ with $a$. By leveraging VLM itself as policy model, $\pi_{\theta}$, we sample candidate steps from the probability distribution $\pi_{\theta}(a|x,I,s_{t})$, where $x$ denotes the task’s input prompt and $I$ represents the input image. The MCTS process starts from the root node, $s_{0}$, representing the beginning of a sentence. It then iteratively proceeds through three key phases—selection, expansion and simulation—which are described in detail in the subsequent paragraphs. In contrast to previous studies, during the data filtering stage with MCTS, we prioritize computational efficiency and comprehensive exploration of the solution space, with our focus centered on self-rewarding setting. Consequently, throughout the MCTS process, we do not employ any pretrained or separately trained process reward models, thereby simplifying and accelerating the procedure. The prompt used for MCTS is shown in Appendix A Table 6.
@@
After selecting a node , we directly utilize the policy $\pi_{\theta}$ to generate several reasoning steps until a final answer is produced or a preset reasoning step limit is reached. Subsequently, we employ the corresponding LLM (in our experiments, the Qwen2.5-VL-7B-Instruct and Qwen2.5-VL-72B-Instruct are used, with Qwen2.5-7B-Instruct serving as the critic model) to compare the generated final answer with the ground truth answer, thereby determining the correctness of the response. If the answer is correct, the MCTS process is terminated and the current iteration number $K$ is recorded; if the answer is incorrect, the visit count $N$ of the selected node is updated and the next iteration commences. Appendix A Table 7 illustrates the prompt employed for the critic model.
@@
We apply this MCTS procedure to the entire collection of 70k data samples and record the iteration number $K$ required to solve each problem, using Qwen2.5-VL-7B-Instruct and Qwen2.5-VL-72B-Instruct as the policy model. In this process, $K$ served as a metric for assessing the difficulty of each sample: a higher $K$ indicates that the VLM requires more extensive exploration to arrive at the correct answer, thereby reflecting a greater level of challenge. Ultimately, we select all samples with $K$ greater than 5, as well as those that remained unsolved after 50 iterations, resulting in a final training set of 11k samples with 7B model and 7.5k samples with 72B model. The data difficulty distribution of 11k training set of 7B model is shown in Figure 4 as an example.